# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

### v1

@@ Abstract.
With the development of large language models (LLMs), the ability to handle longer contexts has become a key capability for Web applications such as cross-document understanding and LLM-powered search systems. However, this progress faces two major challenges: performance degradation due to sequence lengths out-of-distribution, and excessively long inference times caused by the quadratic computational complexity of attention. These issues hinder the application of LLMs in long-context scenarios. In this paper, we propose Dynamic Token-Level KV Cache Selection (TokenSelect), a model-agnostic, training-free method for efficient and accurate long-context inference. TokenSelect builds upon the observation of non-contiguous attention sparsity, using Query-Key dot products to measure per-head KV Cache criticality at token-level. By per-head soft voting mechanism, TokenSelect selectively involves a small number of critical KV cache tokens in the attention calculation without sacrificing accuracy. To further accelerate TokenSelect, we designed the Selection Cache based on observations of consecutive Query similarity and implemented efficient dot product kernel, significantly reducing the overhead of token selection. A comprehensive evaluation of TokenSelect demonstrates up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency, while providing superior performance compared to state-of-the-art long-context inference methods.
Efficient LLMs Inference, Long-context LLMs, Sparse Attention
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM on Web Conference; April 28–May 02, 2025; Sydney, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generation
1. Introduction
@@
With the rapid development of large language models (LLMs), the number of parameters is no longer the sole factor significantly affecting model performance. The ability to effectively process longer context information has become one of the key metrics for evaluating LLMs’ capabilities. The latest Web applications such as cross-document understanding (Bai et al., 2024), LLM-powered search systems (Sharma et al., 2024), repository-level code completion (Zhang et al., 2023; Di et al., 2024), and complex reasoning (OpenAI, [n. d.]) have all placed higher demands on the long-context abilities of LLMs. There are two main difficulties in using pre-trained LLMs for long-context inference. On one hand, LLMs are limited by their context length during pre-training (e.g. Llama 3 only has 8192 tokens). Directly inferencing on longer sequences can lead to severe performance degradation due to reasons including sequence lengths out-of-distribution (Xiao et al., 2024; Han et al., 2024). On the other hand, even if LLMs possess sufficiently large context lengths, the quadratic computational complexity of attention with respect to sequence length makes the response time for long-context inference unbearable.
@@
Figure 1. Distribution of tokens participating in attention computation under different sparsity patterns (indicated by blue dots). TokenSelect can more accurately select critical tokens for attention computation.
@@
Previous works have made numerous attempts to address these difficulties. To extend the context length of LLMs, the current common practice is to perform post-training on long texts (Team et al., 2024; Yang et al., 2024; GLM et al., 2024). However, this approach comes with significant computational costs, particularly in two aspects: the synthesis of high-quality long-text data and the training process on extended sequences. To accelerate long-context inference, many studies focus on the sparsity of attention, attempting to reduce the scale of KV Cache involved in computation. The key to this type of method lies in designing sparse patterns for attention, which can be mainly divided into two categories: one uses predefined sparse patterns (Wang et al., 2019; Zaheer et al., 2020; Xiao et al., 2024; Han et al., 2024), while the other estimates the potential importance of KV Cache during the inference process (Zhang et al., 2024; Oren et al., 2024; Xiao et al., 2024; Tang et al., 2024; Jiang et al., 2024), attempting to select relevant KV Cache tokens into attention calculations. However, the design of these sparse patterns is often heuristically based on historical criticality or coarse-grained criticality estimation of tokens, making it difficult to ensure that the selected tokens are truly critical, thus resulting in sub-optimal performance, as shown in Fig. 1.
@@
In this paper, we further observe the non-contiguous sparsity of attention, revealing the importance of designing more fine-grained dynamic sparse patterns. To this end, we propose TokenSelect, a model-agnostic and training-free approach that utilizes token-level selective sparse attention for efficient long-context inference and length extrapolation. Specifically, for each Query, TokenSelect dynamically calculates token-level per-head criticality for the past KV Cache and selects the $k$ most critical tokens through our head soft vote mechanism, involving them in the attention calculation. This reduces the scale of attention calculation to a constant length familiar to the model, while maintaining almost all of the long-context information, thereby simultaneously addressing the two main difficulties for long-context inference. To reduce the overhead of token selection, TokenSelect manages the KV Cache in token-level pages (Zheng et al., 2024) and design efficient kernel for token selection based on Paged KV Cache management through Triton (Tillet et al., 2019). Furthermore, based on our observation of high similarity between consecutive queries, we have designed the Selection Cache, which allows consecutive similar queries to share token selection results, thereby reducing the selection frequency while ensuring its effectiveness.
@@
We evaluate the performance and efficiency of TokenSelect on three representative long-context benchmarks (Zhang et al., 2024; Hsieh et al., 2024; Bai et al., 2024) using three open-source LLMs (Yang et al., 2024; Dubey et al., 2024; AI et al., 2024). The experimental results demonstrate that our TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation compared to FlashInfer (flashinfer ai, [n. d.]), and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency compared to state-of-the-art long-context inference method (Xiao et al., 2024). Simultaneously, it provides superior performance on three long-text benchmarks. In summary, we make the following contributions:
@@
  • An observation on the non-contiguous sparsity of attention that highlights the importance of token-level selection.
@@
  • TokenSelect, a model-agnostic and training-free method that achieves accurate and efficient long-context inference and length extrapolation, which is compatible with mainstream LLM serving systems and ready for Web applications.
@@
  • A comprehensive evaluation of TokenSelect, demonstrating up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency while exhibiting superior performance.
2. Related Works
@@
Due to computational complexity constraints, current LLMs based on Transformers often utilize limited context lengths during pre-training (Touvron et al., 2023; Dubey et al., 2024; Jiang et al., 2023; Yang et al., 2024; GLM et al., 2024; AI et al., 2024). To extend the long-context capabilities of LLMs, current methods can be broadly categorized into three approaches (Huang et al., 2024; Zhou et al., 2024; Zhao et al., 2024): 1) Modifying positional encodings: A widely adopted method is positional interpolation (Chen et al., 2023).  Chen et al. first proposed linear scaling of RoPE (Su et al., 2024) to map longer positional ranges within the original training window. Subsequent works (bloc97, 2023; emozilla, 2023) further improved this method using Neural Tangent Kernel (NTK) theory (Jacot et al., 2018), achieving longer context windows while maintaining model performance. Methods like YaRN (Peng et al., 2024) and Giraffe (Pal et al., 2023) optimize interpolation effects by adjusting frequency components or introducing temperature parameters. 2) Long-context post-training: This approach extends the model’s context length through additional training steps on longer documents after pre-training (Yang et al., 2024; Tian et al., 2024). It has been widely adopted by leading LLMs (Team et al., 2024; Yang et al., 2024; GLM et al., 2024) with the support of sequence parallelism techniques (Shoeybi et al., 2020; Jacobs et al., 2023; Liu et al., 2024). 3) Incorporating additional memory modules: Notable examples include Transformer-XL (Dai* et al., 2019), Compressive Transformer (Rae et al., 2020), RMT (Bulatov et al., 2022) and Infini-attention (Munkhdalai et al., 2024). Although these methods have expanded the context length of LLMs, long-context inference still faces the challenge of high computational costs.
@@
In state-of-the-art LLMs serving systems (Kwon et al., 2023; Huggingface, 2024; NVIDIA, 2024; Zheng et al., 2024), technologies such as Flash Attention (Dao et al., 2022; Dao, 2024) and Paged Attention (Kwon et al., 2023) have greatly optimized LLMs inference efficiency by improving GPU I/O bottlenecks. However, in long-context inference scenarios, the quadratic computational complexity of attention with respect to sequence length poses new challenges for LLMs inference. Numerous studies focus on the sparsity of attention, selecting partial KV Cache for attention calculations to improve long-context inference efficiency. Sliding window (Wang et al., 2019; Zaheer et al., 2020) is one of the most widely used sparse patterns, reducing complexity to linear by executing attention computations within localized windows. Recent works like StreamingLLM (Xiao et al., 2024) and LM-infinite (Han et al., 2024) retain the initial tokens of the sequence in addition to sliding windows, effectively maintaining LLMs’ performance when processing long sequences. While these approaches are simple to implement, they cannot retain information from long contexts. Another approach focuses on dynamic KV Cache selection during inference. Methods like H2O (Zhang et al., 2024), TOVA (Oren et al., 2024), FastGen (Ge et al., 2024), Scissorhands (Liu et al., 2023), and SnapKV (Li et al., 2024) evaluate token criticality based on historical attention scores, selecting tokens within a limited budget. However, these methods permanently discard parts of the KV Cache, causing information loss from long contexts. To address this, InfLLM (Xiao et al., 2024) introduces Block Memory Units for KV Cache management, retrieving information from long contexts and offloading less-used blocks to CPU. Similarly, QUEST (Tang et al., 2024) proposes query-aware sparsity at page granularity, while MInference (Jiang et al., 2024) optimizes long-context inference using three sparse patterns. Apart from considering all attention heads, some other works (Ribar et al., 2024; Lee et al., 2024; Tang et al., 2024) attempt to focus on only a subset of attention heads. Beyond selection, some other research focuses on KV Cache quantization (Liu et al., 2024; Yang et al., 2024; He et al., 2024; Kang et al., 2024) and merging (Liu et al., 2024; Wan et al., 2024; Zhang et al., 2024; Wang et al., 2024). However, existing methods struggle to be applied in real-world Web applications, both in terms of accuracy and efficiency.
3. Preliminaries
@@
In this section, we first introduce the inference process of LLMs, and then define the Selective Sparse Attention Problem.
3.1. LLMs Inference
@@
Nowadays, mainstream LLMs are primarily based on the Decoder-only Transformer architecture, consisting sequentially of a word embedding layer, a series of transformer layers, and a token prediction head. Each transformer layer includes a multi-head attention (MHA) module and a feed-forward networks (FFN) module. The inference process of LLMs can be divided into two stages: the Prefill Stage and the Decode Stage.
@@
The Prefill Stage is the preparatory phase of the inference process. In this stage, the user’s input is processed layer by layer through a single forward pass of LLMs, generating KV Cache for each layer. The generation of KV Cache is completed by the MHA module. Assuming $\mathbf{X}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is the input of a transformer layer, where $n_{\text{in}}$ is the number of tokens in user’s input sequence and $d$ is the hidden size. The computation of MHA in the Prefill Stage is as follows (simplified to single-head mode):
(1) $[\mathbf{Q}_{\text{prefill}},\mathbf{K}_{\text{prefill}},\mathbf{V}_{\text{prefill}}]=\mathbf{X}_{\text{prefill}}\cdot\left[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\right],$
@@
(2) $\mathbf{O}_{\text{prefill}}=\text{softmax}\left(\frac{\mathbf{Q}_{\text{prefill}}\cdot{\mathbf{K}_{\text{prefill}}}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{prefill}},$
@@
(3) $\mathbf{K}_{\text{decode}}$ $=\left[\mathbf{K}_{\text{cache}},\ \mathbf{X}_{\text{decode}}\cdot\mathbf{W}_{k}\right],\;\mathbf{K}_{\text{cache}}\leftarrow\mathbf{K}_{\text{decode}},$
@@
3.2. Selective Sparse Attention
As discussed in the Sec. 1, the high attention sparsity in LLMs suggests sparse attention as a promising solution for long-context inference challenges. Sparse attention can keep the number of tokens participating in attention computations at a constant scale, rather than increasing with sequence length. Given that predefined sparse patterns are detrimental to performance, we aim to dynamically select crucial tokens for attention computation at each step during the inference process. Therefore, we formalize this problem according to the following definition.

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has driven growing demand for processing extended context sequences in contemporary applications. However, this progress faces two major challenges: performance degradation due to sequence lengths out-of-distribution, and excessively long inference times caused by the quadratic computational complexity of attention. These issues hinder the application of LLMs in long-context scenarios. In this paper, we propose Dynamic Token-Level KV Cache Selection (TokenSelect), a training-free method for efficient and accurate long-context inference. TokenSelect builds upon the observation of non-contiguous attention sparsity, using Query-Key dot products to measure per-head KV Cache criticality at token-level. By per-head soft voting mechanism, TokenSelect selectively involves a few critical KV cache tokens in attention calculation without sacrificing accuracy. To further accelerate TokenSelect, we design the Selection Cache based on observations of consecutive Query similarity and implemented efficient dot product kernel, significantly reducing the overhead. A comprehensive evaluation of TokenSelect demonstrates up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency, while providing superior performance compared to state-of-the-art long-context inference methods.
TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection
Wei Wu1$\dagger$,$\ddagger$, Zhuoshi Pan2$\dagger$,$\ddagger$, Chao Wang1, Liyi Chen1$\ddagger$, Yunchu Bai1, Tianfu Wang1, Kun Fu3, Zheng Wang3,*, Hui Xiong4,5,* 1University of Science and Technology of China, 2Tsinghua University, 3Alibaba Cloud Computing, 4Hong Kong University of Science and Technology (Guangzhou), 5Hong Kong University of Science and Technology
11footnotetext: Corresponding Authors.22footnotetext: Equal Contribution.33footnotetext: Work during internship at Alibaba Cloud Computing.44footnotetext: Code: https://github.com/pzs19/TokenSelect
1 Introduction
@@
With the rapid development of large language models (LLMs), the number of parameters is no longer the sole factor significantly affecting model performance. The ability to effectively process longer context information has become one of the key metrics for evaluating LLMs’ capabilities. The latest applications such as cross-document understanding Bai et al. (2024), LLM-powered search systems Sharma et al. (2024), and complex reasoning OpenAI have all placed higher demands on the long-context abilities of LLMs. There are two main difficulties in using pre-trained LLMs for long-context inference. On one hand, LLMs are limited by their context length during pre-training (e.g. Llama 3 only has 8192 tokens). Directly inferencing on longer sequences can lead to severe performance degradation due to reasons including sequence lengths out-of-distribution Xiao et al. (2024b); Han et al. (2024). On the other hand, even if LLMs possess sufficiently large context lengths, the quadratic computational complexity of attention with respect to sequence length makes the response time for long-context inference unbearable.
@@
Figure 1: Distribution of tokens participating in attention computation under different sparsity patterns (indicated by blue dots). TokenSelect can more accurately select critical tokens for attention computation.
@@
Previous works have made numerous attempts to address these difficulties. To extend the context length of LLMs, the current common practice is to perform post-training on long texts Team et al. (2024); Yang et al. (2024a); GLM et al. (2024). However, this approach comes with significant computational costs, particularly in two aspects: the synthesis of high-quality long-text data and the training process on extended sequences. To accelerate long-context inference, many studies focus on the sparsity of attention, attempting to reduce the scale of KV Cache involved in computation. The key to this type of method lies in designing sparse patterns for attention, which can be mainly divided into two categories: one uses predefined sparse patterns Wang et al. (2019); Zaheer et al. (2020); Xiao et al. (2024b); Han et al. (2024), while the other estimates the potential importance of KV Cache during the inference process Zhang et al. (2024c); Oren et al. (2024); Xiao et al. (2024a); Tang et al. (2024b); Jiang et al. (2024), attempting to select relevant KV Cache tokens into attention calculations. However, the design of these sparse patterns is often heuristically based on historical criticality or coarse-grained criticality estimation of tokens, making it difficult to ensure that the selected tokens are truly critical, thus resulting in sub-optimal performance, as shown in Fig. 1.
@@
In this paper, we further observe the non-contiguous sparsity of attention, revealing the importance of designing more fine-grained dynamic sparse patterns. To this end, we propose TokenSelect, a training-free approach that utilizes token-level selective sparse attention for efficient long-context inference and length extrapolation. Specifically, for each Query, TokenSelect dynamically calculates token-level per-head criticality for the past KV Cache and selects the $k$ most critical tokens through our head soft vote mechanism, involving them in the attention calculation. This reduces the scale of attention calculation to a constant length familiar to the model, while maintaining almost all of the long-context information, thereby simultaneously addressing the two main difficulties for long-context inference. To reduce the overhead of token selection, TokenSelect manages the KV Cache in token-level pages Zheng et al. (2024) and designs efficient kernel for token selection based on paged KV Cache management through Triton Tillet et al. (2019). Furthermore, based on our observation of high similarity between consecutive queries, we have designed the Selection Cache, which allows consecutive similar queries to share token selection results, thereby reducing the selection frequency while ensuring its effectiveness.
@@
We evaluate the performance and efficiency of TokenSelect on three representative long-context benchmarks using three open-source LLMs. The experimental results demonstrate that our TokenSelect can achieve up to 23.84×23.84$\times$23.84 × speedup in attention computation compared to FlashInfer flashinfer ai , and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end inference latency compared to state-of-the-art long-context inference method Xiao et al. (2024a). Simultaneously, it provides superior performance on three long-text benchmarks. In summary, we make the following contributions:
@@
  • An observation on the non-contiguous sparsity of attention that highlights the importance of token-level KV Cache selection.
@@
  • TokenSelect, a training-free method that achieves accurate and efficient long-context inference and length extrapolation, compatible with mainstream LLM serving systems.
@@
  • Comprehensive evaluations of our method, showing up to 23.84×23.84$\times$23.84 × speedup in attention computation and up to 2.28×2.28$\times$2.28 × acceleration in end-to-end latency while exhibiting superior long-context performance.
2 Related Works
@@
Due to computational complexity constraints, current LLMs based on Transformers often utilize limited context lengths during pre-training Touvron et al. (2023); Grattafiori et al. (2024); Jiang et al. (2023); Yang et al. (2024a); GLM et al. (2024); AI et al. (2024). To extend the long-context capabilities of LLMs, current methods can be broadly categorized into three approaches Huang et al. (2024); Zhou et al. (2024); Zhao et al. (2024): 1) Modifying positional encodings: A widely adopted method is positional interpolation Chen et al. (2023).  Chen et al. first proposed linear scaling of RoPE Su et al. (2024) to map longer positional ranges within the original training window. Subsequent works bloc97 (2023); emozilla (2023) further improved this method using Neural Tangent Kernel (NTK) theory Jacot et al. (2018), achieving longer context windows while maintaining model performance. Methods like YaRN Peng et al. (2024) and Giraffe Pal et al. (2023) optimize interpolation effects by adjusting frequency components or introducing temperature parameters. 2) Long-context post-training: This approach extends the model’s context length through additional training steps on longer documents after pre-training Yang et al. (2024c); Tian et al. (2024). It has been widely adopted by leading LLMs Team et al. (2024); Yang et al. (2024a); GLM et al. (2024) with the support of sequence parallelism techniques Shoeybi et al. (2020); Jacobs et al. (2023); Liu et al. (2024b). 3) Incorporating additional memory modules: Notable examples include Transformer-XL Dai* et al. (2019), Compressive Transformer Rae et al. (2020), RMT Bulatov et al. (2022) and Infini-attention Munkhdalai et al. (2024). Although these methods have expanded the context length of LLMs, long-context inference still faces the challenge of high computational costs.
@@
In state-of-the-art LLMs serving systems Kwon et al. (2023); Huggingface (2024); NVIDIA (2024); Zheng et al. (2024), technologies such as Flash Attention Dao et al. (2022); Dao (2024) and Paged Attention Kwon et al. (2023) have greatly optimized LLMs inference efficiency by improving GPU I/O bottlenecks. However, in long-context inference scenarios, the quadratic computational complexity of attention with respect to sequence length poses new challenges for LLMs inference. Numerous studies focus on the sparsity of attention, selecting partial KV Cache for attention calculations to improve long-context inference efficiency. Sliding window Wang et al. (2019); Zaheer et al. (2020) is one of the most widely used sparse patterns, reducing complexity to linear by executing attention computations within localized windows. Recent works like StreamingLLM Xiao et al. (2024b) and LM-infinite Han et al. (2024) retain the initial tokens of the sequence in addition to sliding windows, effectively maintaining LLMs’ performance when processing long sequences. While these approaches are simple to implement, they cannot retain information from long contexts. Another approach focuses on dynamic KV Cache selection during inference. Methods like H2O Zhang et al. (2024c), TOVA Oren et al. (2024), FastGen Ge et al. (2024), Scissorhands Liu et al. (2023), and SnapKV Li et al. (2024) evaluate token criticality based on historical attention scores, selecting tokens within a limited budget. However, these methods permanently discard parts of the KV Cache, causing information loss from long contexts. To address this, InfLLM Xiao et al. (2024a) introduces Block Memory Units for KV Cache management, retrieving information from long contexts and offloading less-used blocks to CPU. Similarly, QUEST Tang et al. (2024b) proposes query-aware sparsity at page granularity, while MInference Jiang et al. (2024) optimizes long-context inference using three sparse patterns. Apart from considering all attention heads, some other works Ribar et al. (2024); Lee et al. (2024); Tang et al. (2024a) attempt to focus on only a subset of attention heads. Beyond selection, some other research focuses on KV Cache quantization Liu et al. (2024c); Yang et al. (2024b); He et al. (2024); Kang et al. (2024) and merging Liu et al. (2024a); Wan et al. (2024); Zhang et al. (2024b); Wang et al. (2024). While existing methods have shown progress, opportunities for further improvement remain in achieving optimal accuracy and computational efficiency for real-world deployment.
3 Preliminaries
@@
In this section, we introduce the inference of LLMs and define the Selective Sparse Attention Problem.
3.1 LLMs Inference
@@
Nowadays, mainstream LLMs are primarily based on the Decoder-only Transformer architecture. Each transformer layer includes a multi-head attention (MHA) and a feed-forward network (FFN). The inference process of LLMs can be divided into two stages: the Prefill Stage and the Decode Stage.
@@
The Prefill Stage is the preparatory phase of the inference process. In this stage, the user’s input is processed layer by layer through a single forward pass of LLMs, generating KV Cache for each layer. The generation of KV Cache is completed by the MHA module. Assuming $\mathbf{X}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is the input of a transformer layer, where $n_{\text{in}}$ is the number of tokens in user’s input sequence and $d$ is the hidden size. The MHA computation in the Prefill Stage is as follows (simplified to single head):
$[\mathbf{Q}_{\text{prefill}},\mathbf{K}_{\text{prefill}},\mathbf{V}_{\text{prefill}}]=\mathbf{X}_{\text{prefill}}\cdot\left[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\right],$ (1)
@@
$\mathbf{O}_{\text{prefill}}=\text{softmax}\left(\frac{\mathbf{Q}_{\text{prefill}}\cdot{\mathbf{K}_{\text{prefill}}}^{\top}}{\sqrt{d}}\right)\cdot\mathbf{V}_{\text{prefill}},$ (2)
@@
$\mathbf{K}_{\text{decode}}$ $=\left[\mathbf{K}_{\text{cache}},\ \mathbf{X}_{\text{decode}}\cdot\mathbf{W}_{k}\right],\;\mathbf{K}_{\text{cache}}\leftarrow\mathbf{K}_{\text{decode}},$ (3)
@@
3.2 Selective Sparse Attention
As discussed in the Sec. 1, the high attention sparsity in LLMs suggests sparse attention as a promising solution for long-context inference challenges. Sparse attention can keep the number of tokens participating in attention computations at a constant scale, rather than increasing with sequence length. Given that predefined sparse patterns are detrimental to performance, we aim to dynamically select crucial tokens at each step during the inference process. Therefore, we formalize this problem according to the following definition.
Definition 1 (Selective Sparse Attention Problem, informal). For current input of length $C$ ($C=1$ in the Decode Stage) and KV Cache of length $N$, assuming there are $H$ attention heads with size of $d_{h}$, let $\mathbf{O}$ be the output of the SDPA:
$\mathbf{O}=\left[\sigma\left(\frac{\mathbf{Q}^{h}\cdot{\left[\mathbf{K}^{h}_{\text{cache}},\ \mathbf{K}^{h}_{\text{current}}\right]}^{\top}}{\sqrt{d}}\right)\cdot[\mathbf{V}^{h}_{\text{cache}},\ \mathbf{V}^{h}_{\text{current}}]\right]_{h=1}^{H},$
where $\sigma$ denotes softmax, $\mathbf{Q}^{h},\mathbf{K}^{h}_{\text{current}},\mathbf{V}^{h}_{\text{current}}\in\mathbb{R}^{C\times d_{h}}$ are Query, Key, Value matrices of current input for head $h$ and $\mathbf{K}^{h}_{\text{cache}},\mathbf{V}^{h}_{\text{cache}}\in\mathbb{R}^{N\times d_{h}}$ represent the KV Cache. Let $\hat{\mathbf{O}}$ be the output of the Selective Sparse Attention:
$\hat{\mathbf{O}}=\left[\sigma\left(\frac{\mathbf{Q}^{h}\cdot{\left[\mathbf{K}^{h}_{\text{select}},\ \mathbf{K}^{h}_{\text{current}}\right]}^{\top}}{\sqrt{d}}\right)\cdot[\mathbf{V}^{h}_{\text{select}},\ \mathbf{V}^{h}_{\text{current}}]\right]_{h=1}^{H},$
$\mathcal{S}\left(\mathbf{Q},\ \mathbf{K}_{\text{cache}}\right)=\mathcal{I},\;\text{where}\;\mathcal{I}\in\mathcal{P}(\{1,\cdots,N\}),$ (6)
$\min_{S}\left\|\mathbf{O}-\hat{\mathbf{O}}\right\|_{2}^{2}.$ (7)
Existing works on long-context inference can be categorized under the Selective Sparse Attention Problem, with variations in the design of the selection function $\mathcal{S}$. Zaheer et al. (2020); Xiao et al. (2024b) have developed input-independent selection functions $\mathcal{S}()$, while Zhang et al. (2024c); Oren et al. (2024); Li et al. (2024) propose Query-independent functions $\mathcal{S}(\mathbf{K}_{\text{cache}})$ for improved performance. Current state-of-the-art methods Xiao et al. (2024a); Tang et al. (2024b); Jiang et al. (2024) utilize Query-aware selection functions $\mathcal{S}(\mathbf{Q},\mathbf{K}_{\text{cache}})$. However, these approaches typically operate at a block-level, which limits their effectiveness and overall performance.
4 Motivations and Observations